# Monotonicity of Recurrence in Random Walks

### 06 Mar 2024 v1

@@ 1. Introduction
For two random walks $X$ and $Y$ on $\mathbb{N}^{2}$, say $X\preceq Y$, or equivalently $Y\succeq X$, if for each state, the probability of going down is weakly greater in $X$ than in $Y$, the probability of going left is weakly greater in $X$ than in $Y$, the probability of going up is weakly greater in $Y$ than in $X$, and the probability of going right is weakly greater in $Y$ than in $X$.
@@
On the other hand, we do find conditions under which monotonicity preserves recurrence.
@@
$\begin{cases}\text{right}&\text{w.p. }r_{o}\\ \text{up}&\text{w.p. }u_{o},\end{cases}\qquad\qquad\begin{cases}\text{left}&\text{w.p. }\ell_{x}\\ \text{right}&\text{w.p. }r_{x}\\ \text{up}&\text{w.p. }u_{x}\end{cases}\qquad\qquad\begin{cases}\text{down}&\text{w.p. }d_{y}\\ \text{up}&\text{w.p. }u_{y}\\ \text{right}&\text{w.p. }r_{y},\end{cases}\qquad\qquad\begin{cases}\text{left}&\text{w.p. }\ell_{q}\\ \text{right}&\text{w.p. }r_{q}\\ \text{up}&\text{w.p. }u_{q}\\ \text{down}&\text{w.p. }d_{q},\end{cases}$
@@
Using Theorems 1.5 and 1.6 and their contrapositives, we obtain the following result that allows for application of monotonicity when neither random walk is inward-homogeneous, though an inward-homogeneous random walk must exist between the two random walks.
@@
Rayleigh’s monotonicity law, introduced by Rayleigh [3] in 1899, may be able to make some statements on this question of when monotonicity of recurrence occurs, but is severely limited by the requirement that both $X$ and $Y$ must be Markov chains that correspond to electric networks (see [1] for a comprehensive overview of Rayleigh’s monotonicity law). A Markov chain corresponds to an electric network if there exist constants $c_{xy}\geq 0$ for every unordered pair of two distinct states $x$ and $y$, such that letting $c_{x}=\sum_{y}c_{xy}$ for all states $x$, assuming only finitely many terms in each sum are positive, we have the transition probability from $x$ to $y$ is $c_{xy}/c_{x}$ and the transition probability from $y$ to $x$ is $c_{xy}/c_{y}$. When the Markov chain is positive recurrent, this is equivalent to the Markov chain being reversible, where the stationary distribution $\pi$ is given by $\pi_{x}\propto c_{x}$. This imposes certain conditions on a random walk on $\mathbb{N}^{2}$: for example, for 4 states $w$, $x$, $y$, and $z$, we have $\frac{c_{wx}}{c_{xy}}\cdot\frac{c_{xy}}{c_{yz}}\cdot\frac{c_{yz}}{c_{wz}}\cdot\frac{c_{wz}}{c_{wx}}=1$. Letting $(w,x,y,z)=((0,0),(1,0),(1,1),(0,1))$ in an inward-homogeneous random walk on $\mathbb{N}^{2}$, this implies $\frac{\ell_{x}d_{q}r_{y}u_{o}}{u_{x}\ell_{q}d_{y}r_{o}}=1$. Considering two other “squares” $((1,0),(2,0),(2,1),(1,1))$ and $((0,1),(1,1),(1,2),(0,2))$ yields two additional conditions $\frac{r_{x}\ell_{q}}{\ell_{x}r_{q}}=1$ and $\frac{u_{q}d_{y}}{u_{y}d_{q}}=1$. These conditions severely limit the applicability of Rayleigh monotonicity to two random walks $X\preceq Y$.
@@ 2.1. Recurrent Random Walk
$\begin{cases}\text{right}&\text{w.p. }\frac{9}{14}\\ \text{up}&\text{w.p. }\frac{5}{14},\end{cases}\qquad\qquad\begin{cases}\text{up}&\text{w.p. }\frac{1}{2}\\ \text{right}&\text{w.p. }\frac{5}{12}\\ \text{left}&\text{w.p. }\frac{1}{12}\end{cases}\qquad\qquad\begin{cases}\text{down}&\text{w.p. }\frac{89}{120}\\ \text{up}&\text{w.p. }\frac{5}{24}\\ \text{right}&\text{w.p. }\frac{1}{20},\end{cases}\qquad\qquad\begin{cases}\text{up}&\text{w.p. }\frac{3}{8}\\ \text{down}&\text{w.p. }\frac{3}{8}\\ \text{left}&\text{w.p. }\frac{5}{24}\\ \text{right}&\text{w.p. }\frac{1}{24},\end{cases}$
@@ 2.2. Transient Random Walk
$\begin{cases}\text{right}&\text{w.p. }\frac{9}{14}\\ \text{up}&\text{w.p. }\frac{5}{14},\end{cases}\qquad\qquad\begin{cases}\text{up}&\text{w.p. }\frac{49}{100}\\ \text{right}&\text{w.p. }\frac{41}{100}\\ \text{left}&\text{w.p. }\frac{1}{10}\end{cases}\qquad\qquad\begin{cases}\text{down}&\text{w.p. }\frac{19}{25}\\ \text{up}&\text{w.p. }\frac{5}{25}\\ \text{right}&\text{w.p. }\frac{1}{25},\end{cases}\qquad\qquad\begin{cases}\text{up}&\text{w.p. }\frac{1}{100}\\ \text{down}&\text{w.p. }\frac{74}{100}\\ \text{left}&\text{w.p. }\frac{21}{100}\\ \text{right}&\text{w.p. }\frac{4}{100},\end{cases}$
@@
To show this Markov chain is transient, consider a random walk $X_{n}$ with $X_{0}=(0,0)$. It suffices to show
@@ 3. Monotonicity in Homogeneous Random Walks
We can now couple $X$ and $Y$ as follows, so that the first two stages of $Y$ coincide with the two stages of $X$. While $X$ and $Y$ may not be at the same state, we first use the same selection of left, right, up, or down based on $\ell_{q}$, $r_{q}$, $u_{q}$, and $d_{q}$. Then $X$ and $Y$ both redirect as necessary, but if $X$ and $Y$ use the same redirection rule, i.e., both are on the positive $x$-axis, both are on the positive $y$-axis, or both are at the origin, the same redirection choice is made. If they use different redirection rules, the choices can be made arbitrarily. Finally, if the selected move for $Y$ is up or right, we inward shift as necessary.
It now remains to verify that this coupling satisfies $Y_{t}\leq X_{t}+s_{t}$ for some $s_{t}\in S$ for all $t$, which we prove by induction. We start with $X_{0}=Y_{0}=(0,0)$, and for fixed $t\geq 1$, suppose that $Y_{t-1}\leq X_{t-1}+s_{t-1}$ for some $s_{t-1}\in S$. Let $Y_{t}^{\prime}$ be the state obtained from $Y_{t}$ by ignoring inward shifting at timestep $t$; it suffices to show $Y_{t}^{\prime}\leq X_{t}+s_{t}$ for some $s_{t}\in S$, as $Y_{t}\leq Y_{t}^{\prime}$ always. The only way $Y_{t}^{\prime}\leq X_{t}+s_{t-1}$ could not hold is if $X$ and $Y$ have different redirection rules (with no redirection being a fourth possibility), as otherwise $Y_{t}^{\prime}-Y_{t-1}=X_{t}-X_{t-1}$ and thus $Y_{t}^{\prime}\leq X_{t}+s_{t-1}$ then follows from the inductive hypothesis.
@@
Case 1: left. Redirections only occur on the $y$-axis. If both $X$ and $Y^{\prime}$ redirect but under different rules, then one of them is at the origin while the other is on the positive $y$-axis. If $Y_{t-1}=(0,0)$, then $Y_{t}^{\prime}\leq(1,1)\leq X_{t}+(1,1)$. Otherwise, $X_{t-1}=(0,0)$ and $Y_{t-1}=(0,2)$, where $Y^{\prime}$ can only move down or right; if $Y^{\prime}$ moves right, then $Y_{t}^{\prime}-X_{t}=(1,1)$ or $Y_{t}^{\prime}-X_{t}=(0,2)$ if $X$ moves up or right, respectively, and if $Y^{\prime}$ moves down, then $Y_{t}^{\prime}\leq X_{t}+(1,1)$.
@@
Case 2: down. The arguments for this case are the same as for case 1, except reflected across the line $y=x$. ∎
@@
The assumptions of irreducibility in Theorem 1.5 can be removed, rephrasing recurrence as recurrence of the state $(0,0)$. The proof of Theorem 1.5 only uses irreducibility to imply ellipticity, which is used to show that if $Y_{t}\in\{(0,0),(2,0),(1,1),(0,2)\}$, then in each of these four cases we have $Y_{t+2}=(0,0)$ with positive probability. However, this follows from the assumptions that $Y\preceq X$ and $X$ is inward-homogeneous. If $Y_{t}=(0,0)$ implies $Y_{t+2}\neq(0,0)$ with probability 1, without loss of generality suppose $Y_{t+1}=(1,0)$ with positive probability. Then $Y$ cannot move left from $(1,0)$, implying $\ell_{q}=\ell_{x}=0$ for $X$, so with some probability $X_{1}=(1,0)$, from which it cannot return to $(0,0)$, contradicting recurrence of $X$. Essentially identical arguments address the cases $Y_{t}=(2,0)$ and $Y_{t}=(0,2)$. For $Y_{t}=(1,1)$, there are four cases for how one could have $Y_{t+2}\neq(0,0)$ with probability 1.
Case 1: $Y$ cannot move down from $(0,1)$ and cannot move left from $(1,0)$. Then $X$ also cannot make these moves, and thus cannot ever return to $(0,0)$, contradicting recurrence of $(0,0)$.
Case 2: $Y$ cannot move down from $(0,1)$ and cannot move down from $(1,1)$. Then $X$ cannot ever move down, yet as $Y_{t}=(1,1)$, with positive probability $X$ can reach $(1,1)$, from which it cannot return to $(0,0)$, contradicting recurrence of $(0,0)$.
@@
Case 4: $Y$ cannot move left from $(1,1)$ and cannot move down from $(1,1)$. Then $X$ cannot move left or down in $\{(i,j):i,j\geq 1\}$, yet as $Y_{t}=(1,1)$, with positive probability $X$ can reach $\{(i,j):i,j\geq 1\}$, from which it cannot return to $(0,0)$, contradicting recurrence of $(0,0)$.
@@
Let $X$ and $Y$ both start at the origin. It suffices to show that we can couple $X$ and $Y$ such that $Y_{t}\geq X_{t}-s_{t}$ for some $s_{t}\in\{(2,0),(1,1),(0,2)\}$ for all $t$. Let $S=\{(2,0),(1,1),(0,2)\}$. This is because $X$ being transient implies that with some positive probability, $X$ visits $\{(i,j):i+j\leq 2\}$ finitely often, and thus $Y$ returns to $(0,0)$ finitely often.
@@
As redirections only occur on the $y$-axis, if both $X$ and $Y^{\prime}$ redirect but under different rules, then one of them is at the orign while the other is on the positive $y$-axis. If $X_{t-1}=(0,0)$, then $Y_{t}^{\prime}\geq(1,1)-(1,1)\geq X_{t}-(1,1)$. Otherwise $Y_{t-1}=(0,0)$, so $X_{t-1}=(0,2)$, where $X$ can only move down or right; if $X$ moves right, then $X_{t}-Y_{t}^{\prime}=(1,1)$ or $X_{t}-Y_{t}^{\prime}=(0,2)$ if $Y^{\prime}$ moves up or right, respectively, and if $X$ moves down, then $Y_{t}^{\prime}\geq X_{t}-(1,1)$.
@@ 4.1. A counterexample to monotonicity
$-k-x_{0}+y_{0}-2$ $\leq\lim_{n\to\infty}\mathbb{E}[S_{n}-S_{1}]=\sum_{n=1}^{\infty}\mathbb{E}[S_{n+1}-S_{n}]\leq\sum_{n=1}^{\infty}-0.0892\cdot\Pr(X_{2n}\not\in\{(0,0),(0,k)\})$
@@ 4.2. Monotonicity in Homogeneous Random Walks
Intuitively, $X\trianglelefteq Y$ means $X$ goes left more than $Y$. Similar to Definition 1.4, we introduce a definition of a homogeneous random walk on the slab such that monotonicity holds.
@@
Our proof is structurally similar to that of Theorem 1.5, so for sake of brevity and clarity we will omit additional elaboration when the reasoning is essentially identical to before. Let $X$ and $Y$ both start at the origin. Let $X_{t}=(i_{t}^{(1)},j_{t}^{(1)})$ and $Y_{t}=(i_{t}^{(2)},j_{t}^{(2)})$. It suffices to show that we can couple $X$ and $Y$ such that $|j_{t}^{(2)}-j_{t}^{(1)}|+\max\{i_{t}^{(2)}-i_{t}^{(1)},0\}\leq 2\lceil k/2\rceil$ for all $t$. We couple $X$ and $Y$ in the same manner as before, where inward shifting only ever changes a right movement in $Y$ to a left movement. We couple the upper corner and boundary redirection rules as described in Definition 4.2, which implies that if one of $X$ and $Y$ is on the upper boundary and the other is at the upper corner, then ignoring inward shifting, if they (both) attempt to move up and the one on the upper boundary redirects right, the one at the upper corner will redirect right as well. We similarly couple the origin and lower boundary redirection rules.
@@
Case 1: $X$ redirects while attempting to go up. Then $Y$ goes up, decreasing the first term by 1 from time $t-1$ to $t$, so the inequality holds regardless of how $X$ redirects, as its movement can only change one term by at most 1.
Case 2: $X$ redirects while attempting to go down. This case is symmetrically equivalent to case 1.
Case 3: $X$ redirects while attempting to go left. Then $Y$ goes left, decreasing the second term by 1, so the inequality holds regardless of how $X$ redirects.
Case 4: $Y$ redirects while attempting to go up. This case follows from the same argument as in case 1.
Case 5: $Y$ redirects while attempting to go down. This case is symmetrically equivalent to case 4.
Case 6: $Y$ redirects while attempting to go left. This implies $i_{t}^{(2^{\prime})}-i_{t}^{(1)}\leq 0$, so we trivially bound $|j_{t}^{(2^{\prime})}-j_{t}^{(1)}|+\max\{i_{t}^{(2^{\prime})}-i_{t}^{(1)},0\}\leq k+0\leq 2\lceil k/2\rceil$.
@@
Case 1: $X$ and $Y$ redirect while attempting to go up. If $X$ is to the right, by 2-periodicity we bound the first term by 1 and the second term by 0. If $Y$ is to the right, meaning $X$ is at the upper corner, the situation is more complicated. If $Y$ redirects down, then (ignoring $X$ moving down as then their movements will be the same) $X$ moves right, so the first term increases by 1 while the second term decreases by 1, so the inequality holds by induction. If $Y$ redirects left then the second term decreases by 1, so the inequality holds regardless of how $X$ redirects. If $Y$ redirects right, our coupling implies $X$ redirects right, which implies the inequality continues to hold.
Case 2: $X$ and $Y$ redirect while attempting to go down. This case is symmetrically equivalent to case 1.
Case 3: $X$ and $Y$ redirect while attempting to go left. As $X$ and $Y$ are both on the $y$-axis, by 2-periodicity they are an even distance apart. We have $\max\{i_{t-1}^{(2)}-i_{t-1}^{(1)},0\}=0$. If $k$ is odd, or if $k$ is even and they are at most $k-2$ apart, then $|j_{t-1}^{(2)}-j_{t-1}^{(1)}|\leq 2\lceil k/2\rceil-2$ and the sum of the two terms can increase by at most 2, yielding the desired inequality. Otherwise, $X$ and $Y$ are on the two corners. If either moves vertically then the first term decreases by 1 so the inequality holds regardless of how the other redirects; otherwise, they both move right, in which case the inequality continues to hold. ∎

### 11 Apr 2024 v2

@@ 1. Introduction
For two random walks $X$ and $Y$ on $\mathbb{N}^{2}$, say $X\preceq Y$ if for each state, the probability of going down is weakly greater in $X$ than in $Y$, the probability of going left is weakly greater in $X$ than in $Y$, the probability of going up is weakly greater in $Y$ than in $X$, and the probability of going right is weakly greater in $Y$ than in $X$.
@@
On the other hand, we find conditions under which recurrence is monotone.
@@
$\begin{cases}\text{right}&\text{w.p. }r_{o}\\ \text{up}&\text{w.p. }u_{o},\end{cases}\qquad\begin{cases}\text{left}&\text{w.p. }\ell_{x}\\ \text{right}&\text{w.p. }r_{x}\\ \text{up}&\text{w.p. }u_{x},\end{cases}\qquad\begin{cases}\text{down}&\text{w.p. }d_{y}\\ \text{up}&\text{w.p. }u_{y}\\ \text{right}&\text{w.p. }r_{y},\end{cases}\qquad\begin{cases}\text{left}&\text{w.p. }\ell_{q}\\ \text{right}&\text{w.p. }r_{q}\\ \text{up}&\text{w.p. }u_{q}\\ \text{down}&\text{w.p. }d_{q},\end{cases}$
@@
Using Theorems 1.5 and 1.6 and their contrapositives, we obtain the following result implying monotonicity when neither random walk is inward-homogeneous, though an inward-homogeneous random walk must exist between the two random walks.
@@
Rayleigh’s monotonicity law, introduced by Rayleigh [3] in 1899, may be able to make some statements on this question of when monotonicity of recurrence occurs, but is severely limited by the requirement that both $X$ and $Y$ must be Markov chains that correspond to electric networks (see [1] for a comprehensive overview). A Markov chain corresponds to an electric network if there exist constants $c_{xy}\geq 0$ for every unordered pair of two distinct states $x$ and $y$, such that letting $c_{x}=\sum_{y}c_{xy}$ for all states $x$, assuming only finitely many terms in each sum are positive, we have the transition probability from $x$ to $y$ is $c_{xy}/c_{x}$ and the transition probability from $y$ to $x$ is $c_{xy}/c_{y}$. When the Markov chain is positive recurrent, this is equivalent to the Markov chain being reversible, where the stationary distribution $\pi$ is given by $\pi_{x}\propto c_{x}$. This imposes certain conditions on a random walk on $\mathbb{N}^{2}$: for example, for 4 states $w$, $x$, $y$, and $z$, we have $\frac{c_{wx}}{c_{xy}}\cdot\frac{c_{xy}}{c_{yz}}\cdot\frac{c_{yz}}{c_{wz}}\cdot\frac{c_{wz}}{c_{wx}}=1$. Letting $(w,x,y,z)=((0,0),(1,0),(1,1),(0,1))$ in an inward-homogeneous random walk on $\mathbb{N}^{2}$, this implies $\frac{\ell_{x}d_{q}r_{y}u_{o}}{u_{x}\ell_{q}d_{y}r_{o}}=1$. Considering two other “squares” $((1,0),(2,0),(2,1),(1,1))$ and $((0,1),(1,1),(1,2),(0,2))$ yields $\frac{r_{x}\ell_{q}}{\ell_{x}r_{q}}=1$ and $\frac{u_{q}d_{y}}{u_{y}d_{q}}=1$. These conditions severely limit the applicability of Rayleigh monotonicity to two random walks $X\preceq Y$.
@@ 2.1. Recurrent Random Walk
$\begin{cases}\text{right}&\text{w.p. }\frac{9}{14}\\ \text{up}&\text{w.p. }\frac{5}{14},\end{cases}\qquad\begin{cases}\text{up}&\text{w.p. }\frac{1}{2}\\ \text{right}&\text{w.p. }\frac{5}{12}\\ \text{left}&\text{w.p. }\frac{1}{12},\end{cases}\qquad\begin{cases}\text{down}&\text{w.p. }\frac{89}{120}\\ \text{up}&\text{w.p. }\frac{5}{24}\\ \text{right}&\text{w.p. }\frac{1}{20},\end{cases}\qquad\begin{cases}\text{up}&\text{w.p. }\frac{3}{8}\\ \text{down}&\text{w.p. }\frac{3}{8}\\ \text{left}&\text{w.p. }\frac{5}{24}\\ \text{right}&\text{w.p. }\frac{1}{24},\end{cases}$
@@ 2.2. Transient Random Walk
$\begin{cases}\text{right}&\text{w.p. }\frac{9}{14}\\ \text{up}&\text{w.p. }\frac{5}{14},\end{cases}\qquad\begin{cases}\text{up}&\text{w.p. }\frac{49}{100}\\ \text{right}&\text{w.p. }\frac{41}{100}\\ \text{left}&\text{w.p. }\frac{1}{10},\end{cases}\qquad\begin{cases}\text{down}&\text{w.p. }\frac{19}{25}\\ \text{up}&\text{w.p. }\frac{5}{25}\\ \text{right}&\text{w.p. }\frac{1}{25},\end{cases}\qquad\begin{cases}\text{up}&\text{w.p. }\frac{1}{100}\\ \text{down}&\text{w.p. }\frac{74}{100}\\ \text{left}&\text{w.p. }\frac{21}{100}\\ \text{right}&\text{w.p. }\frac{4}{100},\end{cases}$
@@
To show transience, consider a random walk $X_{n}$ with $X_{0}=(0,0)$. It suffices to show
@@ 3. Monotonicity in Homogeneous Random Walks
We can now couple $X$ and $Y$ as follows, so that the first two stages of $Y$ coincide with the two stages of $X$. We first use the same selection of left, right, up, or down based on $\ell_{q}$, $r_{q}$, $u_{q}$, and $d_{q}$. Then $X$ and $Y$ both redirect as necessary, but if $X$ and $Y$ use the same redirection rule, i.e., both are on the positive $x$-axis, both are on the positive $y$-axis, or both are at the origin, the same redirection choice is made. Finally, if the selected move for $Y$ is up or right, we inward shift as necessary.
It now remains to verify that this coupling satisfies $Y_{t}\leq X_{t}+s_{t}$ for some $s_{t}\in S$ for all $t$, which we prove by induction. We start with $X_{0}=Y_{0}=(0,0)$, and for fixed $t\geq 1$, suppose that $Y_{t-1}\leq X_{t-1}+s_{t-1}$ for some $s_{t-1}\in S$. Let $Y_{t}^{\prime}\geq Y_{t}$ be the state obtained from $Y_{t}$ by ignoring inward shifting at timestep $t$; it suffices to show $Y_{t}^{\prime}\leq X_{t}+s_{t}$ for some $s_{t}\in S$. The only way $Y_{t}^{\prime}\leq X_{t}+s_{t-1}$ could not hold is if $X$ and $Y$ have different redirection rules (with no redirection being a fourth possibility), as otherwise $Y_{t}^{\prime}-Y_{t-1}=X_{t}-X_{t-1}$ and thus $Y_{t}^{\prime}\leq X_{t}+s_{t-1}$ then follows from the inductive hypothesis.
@@
Case 1: left. Redirections only occur on the $y$-axis. If both $X$ and $Y^{\prime}$ redirect but under different rules, then one of them is at the origin while the other is on the positive $y$-axis. If $Y_{t-1}=(0,0)$, then $Y_{t}^{\prime}\leq(1,1)\leq X_{t}+(1,1)$. Otherwise, $X_{t-1}=(0,0)$ and $Y_{t-1}=(0,2)$, where $Y^{\prime}$ can only move down or right; if $Y^{\prime}$ moves right, then $Y_{t}^{\prime}-X_{t}=(1,1)$ or $(0,2)$ if $X$ moves up or right, respectively, and if $Y^{\prime}$ moves down, then $Y_{t}^{\prime}\leq X_{t}+(1,1)$.
@@
Case 2: down. The argument is identical to case 1, but reflected across the line $y=x$. ∎
@@
The assumptions of irreducibility in Theorem 1.5 can be removed, rephrasing recurrence as recurrence of the state $(0,0)$. The proof of Theorem 1.5 only uses irreducibility to imply ellipticity, which is used to show that if $Y_{t}\in\{(0,0),(2,0),(1,1),(0,2)\}$, then in each of these four cases we have $Y_{t+2}=(0,0)$ with positive probability. However, this follows from the assumptions that $Y\preceq X$ and $X$ is inward-homogeneous. If $Y_{t}=(0,0)$ implies $Y_{t+2}\neq(0,0)$ with probability 1, without loss of generality suppose $Y_{t+1}=(1,0)$ with positive probability. Then $Y$ cannot move left from $(1,0)$, implying $\ell_{q}=\ell_{x}=0$ for $X$, so with some probability $X_{1}=(1,0)$, from which it cannot return to $(0,0)$, contradicting recurrence of $X$. Essentially identical arguments address the cases $Y_{t}=(2,0)$ and $Y_{t}=(0,2)$. For $Y_{t}=(1,1)$, there are four cases for how one could have $Y_{t+2}\neq(0,0)$ with probability 1; we show all contradict recurrence of $(0,0)$ in $X$.
Case 1: $Y$ cannot move down from $(0,1)$ and cannot move left from $(1,0)$. Then $X$ also cannot make these moves, and thus cannot ever return to $(0,0)$.
Case 2: $Y$ cannot move down from $(0,1)$ and cannot move down from $(1,1)$. Then $X$ cannot ever move down, yet as $Y_{t}=(1,1)$, with positive probability $X$ can reach $(1,1)$, from which it cannot return to $(0,0)$.
@@
Case 4: $Y$ cannot move left from $(1,1)$ and cannot move down from $(1,1)$. Then $X$ cannot move left or down in $\{(i,j):i,j\geq 1\}$, yet as $Y_{t}=(1,1)$, with positive probability $X$ can reach $\{(i,j):i,j\geq 1\}$, from which it cannot return to $(0,0)$.
@@
Let $X$ and $Y$ both start at $(0,0)$. Let $S=\{(2,0),(1,1),(0,2)\}$. It suffices to show that we can couple $X$ and $Y$ such that $Y_{t}\geq X_{t}-s_{t}$ for some $s_{t}\in\{(2,0),(1,1),(0,2)\}$ for all $t$. This is because $X$ being transient implies that with some positive probability, $X$ visits $\{(i,j):i+j\leq 2\}$ finitely often, and thus $Y$ returns to $(0,0)$ finitely often.
@@
As redirections only occur on the $y$-axis, if both $X$ and $Y^{\prime}$ redirect but under different rules, then one of them is at the origin while the other is on the positive $y$-axis. If $X_{t-1}=(0,0)$, then $Y_{t}^{\prime}\geq(1,1)-(1,1)\geq X_{t}-(1,1)$. Otherwise $Y_{t-1}=(0,0)$, so $X_{t-1}=(0,2)$, where $X$ can only move down or right; if $X$ moves right, then $X_{t}-Y_{t}^{\prime}=(1,1)$ or $X_{t}-Y_{t}^{\prime}=(0,2)$ if $Y^{\prime}$ moves up or right, respectively, and if $X$ moves down, then $Y_{t}^{\prime}\geq X_{t}-(1,1)$.
@@ 4.1. A counterexample to monotonicity
$-k-x_{0}+y_{0}-2$ $\leq\lim_{n\to\infty}\mathbb{E}[S_{n}-S_{1}]=\sum_{n=1}^{\infty}\mathbb{E}[S_{n+1}-S_{n}]$
$\leq\sum_{n=1}^{\infty}-0.0892\cdot\Pr(X_{2n}\not\in\{(0,0),(0,k)\})$
@@ 4.2. Monotonicity in Homogeneous Random Walks
Intuitively, $X\trianglelefteq Y$ means $X$ goes left more than $Y$. Similar to Definition 1.4, we introduce a definition of homogeneity on the slab such that monotonicity holds.
@@
Our proof is structurally similar to that of Theorem 1.5, so for sake of brevity and clarity we will omit additional elaboration when the reasoning is essentially identical to before. It suffices to show that we can couple $X$ and $Y$ such that $|j_{t}^{(2)}-j_{t}^{(1)}|+\max\{i_{t}^{(2)}-i_{t}^{(1)},0\}\leq 2\lceil k/2\rceil$ for all $t$. We couple $X$ and $Y$ in the same manner as before, where inward shifting only ever changes a right movement in $Y$ to a left movement. We couple the upper corner and boundary redirection rules as described in Definition 4.2, which implies that if one of $X$ and $Y$ is on the upper boundary and the other is at the upper corner, then ignoring inward shifting, if they (both) attempt to move up and the one on the upper boundary redirects right, the one at the upper corner will redirect right as well. We similarly couple the origin and lower boundary redirection rules.
@@
Case 1: $X$ redirects trying to go up. Then $Y$ goes up, decreasing the first term by 1 from time $t-1$ to $t$, so the inequality holds regardless of how $X$ redirects, as its movement can only change one term by at most 1.
Case 2: $X$ redirects trying to go down. This case is symmetrically equivalent to case 1.
Case 3: $X$ redirects trying to go left. Then $Y$ goes left, decreasing the second term by 1, so the inequality holds regardless of how $X$ redirects.
Case 4: $Y$ redirects trying to go up. This case follows identically to case 1.
Case 5: $Y$ redirects trying to go down. This case is symmetrically equivalent to case 4.
Case 6: $Y$ redirects trying to go left. This implies $i_{t}^{(2^{\prime})}-i_{t}^{(1)}\leq 0$, so we trivially bound $|j_{t}^{(2^{\prime})}-j_{t}^{(1)}|+\max\{i_{t}^{(2^{\prime})}-i_{t}^{(1)},0\}\leq k+0\leq 2\lceil k/2\rceil$.
@@
Case 1: $X$ and $Y$ redirect trying to go up. If $X$ is to the right, by 2-periodicity we bound the first term by 1 and the second term by 0. If $Y$ is to the right, meaning $X$ is at the upper corner, the situation is more complicated. If $Y$ redirects down, then (ignoring $X$ moving down as then their movements will be the same) $X$ moves right, so the first term increases by 1 while the second term decreases by 1, so the inequality holds by induction. If $Y$ redirects left then the second term decreases by 1, so the inequality holds regardless of how $X$ redirects. If $Y$ redirects right, our coupling implies $X$ redirects right, which implies the inequality continues to hold.
Case 2: $X$ and $Y$ redirect trying to go down. This case follows symmetrically to case 1.
Case 3: $X$ and $Y$ redirect trying to go left. As $X$ and $Y$ are both on the $y$-axis, by 2-periodicity they are an even distance apart. We have $\max\{i_{t-1}^{(2)}-i_{t-1}^{(1)},0\}=0$. If $k$ is odd, or if $k$ is even and they are at most $k-2$ apart, then $|j_{t-1}^{(2)}-j_{t-1}^{(1)}|\leq 2\lceil k/2\rceil-2$ and the sum of the two terms can increase by at most 2, yielding the desired inequality. Otherwise, $X$ and $Y$ are on the two corners. If either moves vertically then the first term decreases by 1 so the inequality holds regardless of how the other redirects; otherwise, they both move right, in which case the inequality continues to hold. ∎